# Unexpected Improvements to Expected Improvement for Bayesian Optimization

### v2

@@ 7 Conclusion
While our contributions may not apply verbatim to other classes of acquisition functions, our key insights and strategies do translate and could help with e.g. improving information-based (Henrández-Lobato et al., 2014; Wang and Jegelka, 2017), cost-aware (Snoek et al., 2012; Lee et al., 2020), and other types of acquisition functions that are prone to similar numerical challenges. Further, combining the proposed methods with gradient-aware first-order BO methods Ament and Gomes (2022); De Roos et al. (2021); Eriksson et al. (2018) could lead to particularly effective high-dimensional applications of BO, since the advantages of both methods tend to increase with the dimensionality of the search space. Overall, we hope that our findings will increase awareness in the community for the importance of optimizing acquisition functions well, and in particular, for the care that the involved numerics demand.
@@
The authors thank Frank Hutter for valuable references about prior work on numerically stable computations of analytic EI, David Bindel for insightful conversations about the difficulty of optimizing EI, as well as the anonymous reviewers for their knowledgeable feedback.
@@ A.1 Analytic Expected Improvement
Fortunately, this form exposes the quadratic factor, $\Phi(z)/\phi(z)$ can be computed via standard implementations of the scaled complementary error function erfcx, and $\log\left(1+z\Phi(z)/\phi(z)\right)$, the last term of Eq. (12), can be computed stably with the log1mexp implementation proposed in [Mächler, 2012]:
@@ A.2 Analytical LogEI’s Asymptotics
$\log\left(1+\frac{z\Phi(z)}{\phi(z)}\right)$ $=\log\left(\sum_{k=1}^{K}(-1)^{k+1}\left[\prod_{j=0}^{k-1}(2j+1)\right]z^{-2k}\right)+\mathcal{O}(|z^{-2K}|).$ (16)
@@
We first derived a Laurent expansion of the non-log-transformed $z\Phi(z)/\phi(z)$, a key quantity in the last term of Eq. (12), with the help of Wolfram Alpha [Inc., 2023]:
$\frac{z\Phi(z)}{\phi(z)}=-1-\sum_{k=1}^{K}(-1)^{k}\left[\prod_{j=0}^{k-1}(2j+1)\right]z^{-2k}+\mathcal{O}(|z|^{-2(K+1)}).$ (17)
@@
$\log\left(1+\frac{z\Phi(z)}{\phi(z)}\right)$ $=\log\left(L(z,K)+\mathcal{O}(|z|^{-2(K+1)})\right)$ (18)
$=\log L(z,K)+\log(1+\mathcal{O}(|z|^{-2(K+1)})/L(z,K))$
$=\log L(z,K)+\mathcal{O}(\mathcal{O}(|z|^{-2(K+1)})/L(z,K))$
$=\log L(z,K)+\mathcal{O}(|z|^{-2K}).$
@@
Figure 10: Left: Comparison of LogEI values in single-precision floating point arithmetic, as a function of $z=(\mu(x)-f^{*})/\sigma(x)$ to RoBO’s LogEI implementation [Klein et al., 2017]. Notably, RoBO’s LogEI improves greatly on the naïve implementation (Fig. 8), but still exhibits failure points (red) well above floating point underflow. The implementation of Eq. (14) continues to be stable in this regime. Right: Comparison of LogEI values to HEBO’s approximate LogEI implementation [Cowen-Rivers et al., 2022]. At the time of writing, HEBO’s implementation exhibits a discontinuity and error of $>1.02$ at the threshold $z=-6$, below which the approximation takes effect. The discontinuity could be ameliorated, though not removed, by correcting the log normalization constant (turquoise). The figure also shows that the naïve implementation used by HEBO for $z>-6$ starts to become unstable well before $z=-6$.
@@
To our knowledge, SMAC 1.0’s implementation of the logarithm of analytic EI due to Hutter et al. [2011], later translated to RoBO Klein et al. [2017], was the first to improve the numerical stability of analytic EI through careful numerics. The associated implementation is mathematically identical to $\log\circ\ {\text{EI}}{}$, and greatly improves the numerical stability of the computation. For large negative $z$ however, the implementation still exhibits instabilities that gives rise to floating point infinities through which useful gradients cannot be propagated (Fig. 10, left). The implementation proposed herein remedies this problem by switching to the asymptotic approximation of Eq. (15) once it is accurate to machine precision $\epsilon$. This is similar to the use of asymptotic expansions for the computation of $\alpha$-stable densities proposed by Ament and O’Neil [2018].
@@
HEBO Cowen-Rivers et al. [2022] contains an approximation to the logarithm of analytical EI as part of its implementation of the MACE acquisition function Lyu et al. [2018], which – at the time of writing – is missing the log normalization constant of the Gaussian density, leading to a large discontinuity at the chosen cut-off point of $z=-6$ below which the approximation takes effect, see here for the HEBO implementation. Notably, HEBO does not implement an exact stable computation of LogEI like the one of Hutter et al. [2011], Klein et al. [2017] and the non-asymptotic branches of the current work. Instead, it applies the approximation for all $z<-6$, where the implementation exhibits a maximum error of $>1.02$, or if the implementation’s normalization constant were corrected, a maximum error of $>0.1$. By comparison, the implementation put forth herein is mathematically exact for the non-asymptotic regime $z>-1/\sqrt{\epsilon}$ and accurate to numerical precision in the asymptotic regime due to the design of the threshold value.
@@ A.4 A Class of Smooth Approximations with Fat Tails for Larger Batches
To solve this problem, we propose a new smooth approximation to the ReLU, maximum, and indicator functions that decay only polynomially as $x\to-\infty$, instead of exponentially, like the canonical softplus. The high level idea is to use $(1+x^{2})^{-1}$, which is proportional to the Cauchy density function (and is also known as a Lorentzian), in ways that maintain key properties of existing smooth approximations – convexity, positivity, etc – while changing the asymptotic behavior of the functions from exponential to $\mathcal{O}(1/x^{2})$ as $x\to-\infty$, also known as a “fat tail”. Further, we will show that the proposed smooth approximations satisfy similar maximum error bounds as their exponentially decaying counterparts, thereby permitting a similar approximation guarantee as Lemma 2 with minor adjusments to the involved constants. While the derivations herein are based on the Cauchy density with inverse quadratic decay, it is possible to generalize the derivations to e.g. $\alpha$-stable distribution whose symmetric variants permit accurate and efficient numerical computation Ament and O’Neil [2018].
@@ A.6 Parallel Expected Hypervolume Improvement
The hypervolume improvement can be computed via the inclusion-exclusion principle, see [Daulton et al., 2020] for details, we focus on the numerical issues concerning qEHVI here. To this end, we define
@@
where $\mathbf{f}$ is the vector-valued objective function, and $\mathbf{u}_{k}$ is the vector of upper bounds of one of $K$ hyper-rectangles that partition the non-Pareto-dominated space, see [Daulton et al., 2020] for details on the partitioning. Letting $\mathbf{l}_{k}$ be the corresponding lower bounds of the hyper-rectangles, the hypervolume improvement can then be computed as
@@ A.7 Probability of Improvement
The same numerical improvements used by qLogEI to improve Monte-Carlo expected improvement (qEI) in Appendix A.3 can be applied to improve the fully Monte Carlo Noisy Expected Improvement [Letham et al., 2019, Balandat et al., 2020] acquisition function. As in qLogEI, we can (i) approximate the $\max(0,{\mathbf{x}})$ using a softplus to smooth the sample-level improvements to ensure that they are mathematically positive, (ii) approximate the maximum over the $q$ candidate designs by norm $||\cdot||_{\frac{1}{\tau_{\max}}}$, and (iii) take the logarithm to of the resulting smoothed value to mitigate vanishing gradients. To further mitigate vanishing gradients, we can again leverage the Fat Softplus and Fat Maximum approximations. The only notable difference in the $q$EI and $q$NEI acquisition functions is the choice of incumbent, and similarly only a change of incumbent is required to obtain qLogNEI from qLogEI. Specifically, when the scalar $y^{*}$ in Equation (10) is replaced with a vector containing the new incumbnent under each sample wer obtain the qLogNEI acquisition value. The $i^{\text{th}}$ element of the incumbent vector for qLogNEI is $\max_{j^{\prime}=q+1}^{n+q}\xi^{i}({\mathbf{x}}_{j^{\prime}})$, where ${\mathbf{x}}_{q+1},...,{\mathbf{x}}_{n+q}$ are the previously evaluated designs and $\xi^{i}({\mathbf{x}}_{j^{\prime}})$ is the value of the $j^{\prime\text{th}}$ point under the $i^{\text{th}}$ sample from the joint posterior over ${\mathbf{x}}_{1},...,{\mathbf{x}}_{n+q}$. We note that we use a hard maximum to compute the incumbent for each sample because we do not need to compute gradients with respect to the previously evaluated designs ${\mathbf{x}}_{q+1},...,{\mathbf{x}}_{n+q}$.
We note that we obtain computational speed ups by (i) pruning the set of previously evaluated points that considered for being the best incumbent to include only those designs with non-zero probability of being the best design and (ii) caching the Cholesky decomposition of the posterior covariance over the resulting pruned set of previously evaluated designs and using low-rank updates for efficient sampling [Daulton et al., 2022a].
@@ Appendix B Strategies for Optimizing Acquisition Functions
As discussed in Section 2.3, a variety of different approaches and heuristics have been applied to the problem of optimizing acquisition functions. For the purpose of this work, we only consider continuous domains $\mathbb{X}$. While discrete and/or mixed domains are also relevant in practice and have received substantial attention in recent years – see e.g. Baptista and Poloczek [2018], Oh et al. [2019], Kim et al. [2022], Wan et al. [2021], Daulton et al. [2022b], Deshwal et al. [2023] – our work here on improving acquisition functions is largely orthogonal to this (though the largest gains should be expected when using gradient-based optimizers, as is done in mixed-variable BO when conditioning on discrete variables, or when performing discrete or mixed BO using continuous relaxations, probabilistic reparameterization, or straight-through estimators [Daulton et al., 2022b]).
@@
Another relatively straightforward approach is to use zeroth-order methods such as DIRECT [Jones et al., 1993] (used e.g. by Dragonfly [Kandasamy et al., 2020]) or the popular CMA-ES [Hansen et al., 2003]. These approaches are easy implement as they avoid the need to compute gradients of acquisition functions. However, not relying on gradients is also what renders their optimization performance inferior to gradient based methods, especially for higher-dimensional problems and/or joint batch optimization in parallel Bayesian optimization.
The most common approach to optimizing acquisition functions on continuous domains is using gradient descent-type algorithms. Gradients are either be computed based on analytically derived closed-form expressions, or via auto-differentiation capabilities of modern ML systems such as PyTorch [Paszke et al., 2019], Tensorflow [Abadi et al., 2015], or JAX [Bradbury et al., 2018].
For analytic acquisition functions, a common choice of optimizer is L-BFGS-B [Byrd et al., 1995], a quasi-second order method that uses gradient information to approximate the Hessian and supports box constraints. If other, more general constraints are imposed on the domain, other general purpose nonlinear optimizers such as SLSQP [Kraft, 1988] or IPOPT Wächter and Biegler [2006] are used (e.g. by BoTorch). For Monte Carlo (MC) acquisition functions, Wilson et al. [2018] proposes using stochastic gradient ascent (SGA) based on stochastic gradient estimates obtained via the reparameterization trick [Kingma and Welling, 2013]. Stochastic first-order algorithms are also used by others, including e.g. Wang et al. [2016] and Daulton et al. [2022b]. Balandat et al. [2020] build on the work by Wilson et al. [2018] and show how sample average approximation (SAA) can be employed to obtain deterministic gradient estimates for MC acquisition functions, which has the advantage of being able to leverage the improved convergence rates of optimization algorithms designed for deterministic functions such as L-BFGS-B. This general approach has since been used for a variety of other acquisition functions, including e.g. Daulton et al. [2020] and Jiang et al. [2020].
Very few implementations of Bayesian Optimization actually use higher-order derivative information, as this either requires complex derivations of analytical expressions and their custom implementation, or computation of second-order derivatives via automated differentiation, which is less well supported and computationally much more costly than computing only first-order derivatives. One notable exception is Cornell-MOE [Wu and Frazier, 2016, Wu et al., 2017], which supports Newton’s method (though this is limited to the acquisition functions implemented in C++ within the library and not easily extensible to other acquisition functions).

### v3

@@ 7 Conclusion
While our contributions may not apply verbatim to other classes of acquisition functions, our key insights and strategies do translate and could help with e.g. improving information-based [36, 78], cost-aware [68, 53], and other types of acquisition functions that are prone to similar numerical challenges. Further, combining the proposed methods with gradient-aware first-order BO methods [5, 16, 23] could lead to particularly effective high-dimensional applications of BO, since the advantages of both methods tend to increase with the dimensionality of the search space. Overall, we hope that our findings will increase awareness in the community for the importance of optimizing acquisition functions well, and in particular, for the care that the involved numerics demand.
@@
The authors thank Frank Hutter and Peter Frazier for valuable references about prior work on numerically stable computations of analytic EI, David Bindel for insightful conversations about the difficulty of optimizing EI, as well as the anonymous reviewers for their knowledgeable feedback.
@@ A.1 Analytic Expected Improvement
Fortunately, this form exposes the quadratic factor, $\Phi(z)/\phi(z)$ can be computed via standard implementations of the scaled complementary error function erfcx, and $\log\left(1+z\Phi(z)/\phi(z)\right)$, the last term of Eq. (12), can be computed stably with the log1mexp implementation proposed in [58]:
@@ A.2 Analytical LogEI’s Asymptotics
$\log\left(1+\frac{z\Phi(z)}{\phi(z)}\right)$ $=\log\left(\sum_{k=1}^{K}(-1)^{k+1}\left[\prod_{j=0}^{k-1}(2j+1)\right]z^{-2k}\right)+\mathcal{O}(|z^{-2(K-1)}|).$ (16)
@@
We first derived a Laurent expansion of the non-log-transformed $z\Phi(z)/\phi(z)$, a key quantity in the last term of Eq. (12), with the help of Wolfram Alpha [39]:
$\frac{z\Phi(z)}{\phi(z)}=-1-\sum_{k=1}^{K}(-1)^{k}\left[\prod_{j=0}^{k-1}(2j+1)\right]z^{-2k}+\mathcal{O}(|z|^{-2K}).$ (17)
@@
$\log\left(1+\frac{z\Phi(z)}{\phi(z)}\right)$ $=\log\left(L(z,K)+\mathcal{O}(|z|^{-2K})\right)$ (18)
$=\log L(z,K)+\log(1+\mathcal{O}(|z|^{-2K})/L(z,K))$
$=\log L(z,K)+\mathcal{O}(\mathcal{O}(|z|^{-2K})/L(z,K))$
$=\log L(z,K)+\mathcal{O}(|z|^{-2(K-1)}).$
@@
Figure 10: Left: Comparison of LogEI values in single-precision floating point arithmetic, as a function of $z=(\mu(x)-f^{*})/\sigma(x)$ to RoBO’s LogEI implementation [48]. Notably, RoBO’s LogEI improves greatly on the naïve implementation (Fig. 8), but still exhibits failure points (red) well above floating point underflow. The implementation of Eq. (14) continues to be stable in this regime. Right: Comparison of LogEI values to HEBO’s approximate LogEI implementation [12], as well as via the Mills ratio approximation proposed by Frazier et al. [25]. At the time of writing, HEBO’s implementation exhibits a discontinuity and error of $>1.02$ at the threshold $z=-6$, below which the approximation takes effect. The discontinuity could be ameliorated, though not removed, by correcting the log normalization constant (turquoise). The figure also shows that the baseline implementation used by HEBO for $z>-6$ starts to become unstable well before $z=-6$, and that its asymptotic accuracy is lower the accuracy of the Mills ratio approximation.
Approximate Analytic LogEI via Mills Ratio
Frazier et al. [25, 27] are among the earliest references to propose improvements to the numerical computation of EI in the asymptotic regime. Frazier et al. [25] propose approximating the Mills ratio as $\Phi(z)/\phi(z)\approx z/(z^{2}+1)$ to avoid numerical issues in the computation of the knowledge gradient (KG) for $z<-10$. Based on this, Frazier et al. [27] propose an approximation of EI given by $\phi(z)/(z^{2}+1)$, which can readily be converted to log space. By comparison, the analytic EI formulation proposed herein is mathematically exact for $z>-10^{6}$ by leveraging implementations of the scaled complementary error function erfcx for the computation of the Mills ratio $\Phi(z)/\phi(z)$ to numerical precision, in conjunction with log-space computations as detailed in Section A.1. See Figure 10 (right) for a numerical comparison.
@@
To our knowledge, SMAC 1.0’s implementation of the logarithm of analytic EI due to Hutter et al. [37], later translated to RoBO [48], was among the first to improve the stability of analytic EI through careful numerics without asymptotic approximations. The associated implementation is mathematically identical to $\log\circ\ {\text{EI}}{}$, and greatly improves the numerical stability of the computation. For large negative $z$ however, the implementation still exhibits instabilities that gives rise to floating point infinities through which useful gradients cannot be propagated (Fig. 10, left). The implementation proposed herein remedies this by switching to the asymptotic approximation of Eq. (15) once it is accurate to machine precision $\epsilon$, similar to the use of asymptotic expansions for the computation of $\alpha$-stable densities proposed by Ament and O’Neil [2].
@@
HEBO [12] contains an approximation to the logarithm of analytical EI as part of its implementation of the MACE acquisition function [57], which – at the time of writing – is missing the log normalization constant of the Gaussian density ($\log(2\pi)/2$), leading to a large discontinuity at the chosen cut-off point of $z=-6$ below which the approximation takes effect, see here for the HEBO implementation. Notably, HEBO does not implement an exact stable computation of LogEI like the one of Hutter et al. [37], Klein et al. [48] and the non-asymptotic branches of the current work. Instead, it applies the approximation for all $z<-6$, where the implementation exhibits a maximum error of $>1.02$, or if the implementation’s normalization constant were corrected, a maximum error of $>0.1$. By comparison, the implementation put forth herein is mathematically exact for the non-asymptotic regime $z>-1/\sqrt{\epsilon}$ and accurate to numerical precision in the asymptotic regime due to the design of the threshold value.
@@ A.4 A Class of Smooth Approximations with Fat Tails for Larger Batches
To solve this problem, we propose a new smooth approximation to the ReLU, maximum, and indicator functions that decay only polynomially as $x\to-\infty$, instead of exponentially, like the canonical softplus. The high level idea is to use $(1+x^{2})^{-1}$, which is proportional to the Cauchy density function (and is also known as a Lorentzian), in ways that maintain key properties of existing smooth approximations – convexity, positivity, etc – while changing the asymptotic behavior of the functions from exponential to $\mathcal{O}(1/x^{2})$ as $x\to-\infty$, also known as a “fat tail”. Further, we will show that the proposed smooth approximations satisfy similar maximum error bounds as their exponentially decaying counterparts, thereby permitting a similar approximation guarantee as Lemma 2 with minor adjusments to the involved constants. While the derivations herein are based on the Cauchy density with inverse quadratic decay, it is possible to generalize the derivations to e.g. $\alpha$-stable distribution whose symmetric variants permit accurate and efficient numerical computation [2].
@@ A.6 Parallel Expected Hypervolume Improvement
The hypervolume improvement can be computed via the inclusion-exclusion principle, see [13] for details, we focus on the numerical issues concerning qEHVI here. To this end, we define
@@
where $\mathbf{f}$ is the vector-valued objective function, and $\mathbf{u}_{k}$ is the vector of upper bounds of one of $K$ hyper-rectangles that partition the non-Pareto-dominated space, see [13] for details on the partitioning. Letting $\mathbf{l}_{k}$ be the corresponding lower bounds of the hyper-rectangles, the hypervolume improvement can then be computed as
@@ A.7 Probability of Improvement
The same numerical improvements used by qLogEI to improve Monte-Carlo expected improvement (qEI) in Appendix A.3 can be applied to improve the fully Monte Carlo Noisy Expected Improvement [54, 6] acquisition function. As in qLogEI, we can (i) approximate the $\max(0,{\mathbf{x}})$ using a softplus to smooth the sample-level improvements to ensure that they are mathematically positive, (ii) approximate the maximum over the $q$ candidate designs by norm $||\cdot||_{\frac{1}{\tau_{\max}}}$, and (iii) take the logarithm to of the resulting smoothed value to mitigate vanishing gradients. To further mitigate vanishing gradients, we can again leverage the Fat Softplus and Fat Maximum approximations. The only notable difference in the $q$EI and $q$NEI acquisition functions is the choice of incumbent, and similarly only a change of incumbent is required to obtain qLogNEI from qLogEI. Specifically, when the scalar $y^{*}$ in Equation (10) is replaced with a vector containing the new incumbnent under each sample wer obtain the qLogNEI acquisition value. The $i^{\text{th}}$ element of the incumbent vector for qLogNEI is $\max_{j^{\prime}=q+1}^{n+q}\xi^{i}({\mathbf{x}}_{j^{\prime}})$, where ${\mathbf{x}}_{q+1},...,{\mathbf{x}}_{n+q}$ are the previously evaluated designs and $\xi^{i}({\mathbf{x}}_{j^{\prime}})$ is the value of the $j^{\prime\text{th}}$ point under the $i^{\text{th}}$ sample from the joint posterior over ${\mathbf{x}}_{1},...,{\mathbf{x}}_{n+q}$. We note that we use a hard maximum to compute the incumbent for each sample because we do not need to compute gradients with respect to the previously evaluated designs ${\mathbf{x}}_{q+1},...,{\mathbf{x}}_{n+q}$.
We note that we obtain computational speed ups by (i) pruning the set of previously evaluated points that considered for being the best incumbent to include only those designs with non-zero probability of being the best design and (ii) caching the Cholesky decomposition of the posterior covariance over the resulting pruned set of previously evaluated designs and using low-rank updates for efficient sampling [14].
@@ Appendix B Strategies for Optimizing Acquisition Functions
As discussed in Section 2.3, a variety of different approaches and heuristics have been applied to the problem of optimizing acquisition functions. For the purpose of this work, we only consider continuous domains $\mathbb{X}$. While discrete and/or mixed domains are also relevant in practice and have received substantial attention in recent years – see e.g. Baptista and Poloczek [7], Oh et al. [64], Kim et al. [46], Wan et al. [76], Daulton et al. [15], Deshwal et al. [18] – our work here on improving acquisition functions is largely orthogonal to this (though the largest gains should be expected when using gradient-based optimizers, as is done in mixed-variable BO when conditioning on discrete variables, or when performing discrete or mixed BO using continuous relaxations, probabilistic reparameterization, or straight-through estimators [15]).
@@
Another relatively straightforward approach is to use zeroth-order methods such as DIRECT [43] (used e.g. by Dragonfly [45]) or the popular CMA-ES [34]. These approaches are easy implement as they avoid the need to compute gradients of acquisition functions. However, not relying on gradients is also what renders their optimization performance inferior to gradient based methods, especially for higher-dimensional problems and/or joint batch optimization in parallel Bayesian optimization.
The most common approach to optimizing acquisition functions on continuous domains is using gradient descent-type algorithms. Gradients are either be computed based on analytically derived closed-form expressions, or via auto-differentiation capabilities of modern ML systems such as PyTorch [65], Tensorflow [1], or JAX [9].
For analytic acquisition functions, a common choice of optimizer is L-BFGS-B [10], a quasi-second order method that uses gradient information to approximate the Hessian and supports box constraints. If other, more general constraints are imposed on the domain, other general purpose nonlinear optimizers such as SLSQP [49] or IPOPT [75] are used (e.g. by BoTorch). For Monte Carlo (MC) acquisition functions, Wilson et al. [79] proposes using stochastic gradient ascent (SGA) based on stochastic gradient estimates obtained via the reparameterization trick [47]. Stochastic first-order algorithms are also used by others, including e.g. Wang et al. [77] and Daulton et al. [15]. Balandat et al. [6] build on the work by Wilson et al. [79] and show how sample average approximation (SAA) can be employed to obtain deterministic gradient estimates for MC acquisition functions, which has the advantage of being able to leverage the improved convergence rates of optimization algorithms designed for deterministic functions such as L-BFGS-B. This general approach has since been used for a variety of other acquisition functions, including e.g. Daulton et al. [13] and Jiang et al. [42].
Very few implementations of Bayesian Optimization actually use higher-order derivative information, as this either requires complex derivations of analytical expressions and their custom implementation, or computation of second-order derivatives via automated differentiation, which is less well supported and computationally much more costly than computing only first-order derivatives. One notable exception is Cornell-MOE [80, 81], which supports Newton’s method (though this is limited to the acquisition functions implemented in C++ within the library and not easily extensible to other acquisition functions).